# MetaTool Benchmark for Large Language Models: Deciding Whether to Use Tools and Which to Use

### v5

@@ 1 Introduction
Dimension | APIBank | GPT4Tool | APIBench | ToolLLM | ToolBench | ToolQA | MetaTool |
(Li et al., 2023d) | (Yang et al., 2023c) | (Patil et al., 2023) | (Qin et al., 2023b) | (Xu et al., 2023) | Zhuang et al. (2023) | (Ours) | |
Evaluation Range | ①③④ | ③④ | ③④ | ③④ | ②③④ | ③④ | ①② |
Number of Tasks | 1 | 1 | 1 | 1 | 1 | 1 | 4 |
Reliability Test | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ |
Multi-Tool Test | ✘ | ✘ | ✘ | ✔ | ✘ | ✔ | ✔ |
@@ 2.3.2 Tool Selection
Sub-task 2: tool selection in specific scenarios. The objective of this task is to simulate how LLMs perform using tools when they act as controllers of a system (Shen et al., 2023) faced with different scenarios. As LLMs are widely applied across various domains like biomedical domain (biomedgpt) and educational domain (Kasneci et al., 2023), in scenarios where the system caters to diverse demographics or professions (e.g., software engineer (Qian et al., 2023)), its set of tools also varies. This task allows us to explore the performance disparities of LLMs in selecting different kinds of tools, essentially highlighting a form of bias inherent to LLMs (Ferrara, 2023). In such cases, this task examines how effectively LLMs utilize the tools. Given a query with its label , we specify the tool list containing candidates according to its corresponding scenarios. This task consists of two types of scenarios: the first one is the popularity-related scenario, and the second one is the group-related scenario. For the popularity-related scenario, we have selected the 5, 10, and 15 most popular tools based on the number of tools it’s merged with (refer to Table 9 in the Appendix A.5 for details.) to construct the tool list. As for the group-related scenario, we chose six usual occupations or identities and manually curated a tool list consisting of the 10 most relevant tools for each of them (see Table 10 in the Appendix A.5 for details).
Sub-task 3: tool selection with possible reliability issues. The reliability of LLMs’ tool selection is of utmost importance. However, issues like hallucination (Ji et al., 2023) and sycophancy (Wei et al., 2023) within LLMs’ responses that will negatively impact their selection of tools. Therefore, we introduce sub-task 3. In this task, given a query and its corresponding tool , we need to construct the tool list and ensure . This aims to assess whether LLMs can answer questions honestly and avoid issues like choosing non-existent tools or selecting unrelated tools. It should be noted that this task closely resembles real-world scenarios, as not all existing tools capable of addressing user queries are present in the tool list controlled by LLMs. To be specific, we obtain the embedding of ’s description and get the top-k similar tools about as the way in Task 1. Then we randomly sample tools from the rest tool set to construct , denoted as where and . Overall, we remove the ground-truth tool of query and s most similar tools to keep the tools in not related to as much as possible.
@@ A.4 Flexibility of ToolE
Tool |
Description |
Reason |
airqualityforeast |
Planning something outdoors? Get the 2-day air quality forecast for your city. |
A |
Now |
Get Google Trends. In Japan, you can also get Twitter trends and search Twitter keywords. |
A |
abc_to_audio |
Converts ABC music notation to WAV, MIDI, and PostScript files. |
B |
ChatOCR |
The best way to read text from from any document. ChatOCR will scan and read aloud any text document you provide. |
B |
FinanceTool |
Begin an exciting journey through time, interact with historical events, and learn about the past in a fun and engaging way. |
C |
LawTool |
Enables quick search functionality for relevant laws. |
C |
TicTacToe |
Playing a game of Tic Tac Toe with varying board sizes. You can submit your move and get the AI’s response move. |
D |
Planfit |
Get your tailored workout plan and instructions with videos - AI-powered Workout Coach, Planfit. |
D |
@@ A.5 Others Statistics of ToolE
I can’t seem to remember anything I study. I need help with learning and retaining information effectively. |
Emotion-angry |
I’m feeling really down today, can you summarize this YouTube video for me? [Link to YouTube video] |
Emotion-depressed |
Good day! I’m hoping to find some amazing bargains and discounts today. Can you guide me in the right direction? |
Emotion-happy |
Hey Chatbot, I’m looking for a new pair of sneakers on GoFynd. Can you help me find the latest designs and recommend some popular brands? |
Emotion-excited |
I’m interested in Japanese cosmetics. Can you recommend some good brands? |
Direct-request |
Draw a state diagram for a vending machine. |
Direct-order |
Twitter Trends: What are the trending topics on Twitter in Japan? |
Keyword |
Could you recommend some popular shopping malls or markets in Singapore that offer a wide range of products, including local and international brands, diverse food options, and unique shopping experiences? |
Details |
@@
Software Engineer |
RepoTool, AI2sql, SSH, AutoInfra1, noteable, dart, hackit_web_scanner, LarkBaseImporter, webhooks, universal |
(von Werra & Allal, 2023; Odendahl, 2023; Welsh, 2023; Bridgwater, 2023; Mendelevitch, 2023) |
Elders |
NewsTool, PolishTool, CharityTool, MapTool, MemoryTool, WeatherTool, RestaurantBookingTool, DietTool, NotesTool, TripAdviceTool |
Finance Staff |
FinanceTool, ChartTool, PDF&URLTool, NotesTool, ExchangeTool, CreditYelp, LawTool, DataRetrievalTool, fundsdbsearch, CompanyInfoTool |
(Raieli, 2023; Zhang, 2023; Vidal, 2023; Yang et al., 2023b; Team, 2023) |
House manager |
tira, Discount, ProductSearch, ABCmouse, RestaurantBookingTool, IndoorPlants, recipe_retrieval, HouseRentingTool, TripTool, CreditYelp |
(Empresaria, 2023) |
Students |
CourseTool, ResearchFinder, ResearchHelper, speak, noteable, search, MemoryTool, NotesTool, MixerBox_Translate_AI_language_tutor, ABCmouse |
Artists&designers |
placid, find_agency, ArtCollection, ChartTool, storybird_stories, MediaModifyTool, PolishTool, MusicTool, ImageSearch, BookTool |
@@ C.3 Task Comparison
Tool selection with similar choices |
Tool and its most similar tools. |
Tool selection in specific scenarios |
Specified tools in a certain scenario. |
Tool selection with possible reliability issues |
The tools randomly chosen are from the remaining set of tools, which excludes tool and its most similar tools. |
Multi-tool Selection |
The tools randomly chosen are from the remaining set of tools, which excludes the union set of most similar tools of , where is the ground-truth tool set. |
@@ Appendix E Failure Case Study
1 |
I need to remember to attend a meeting at 10 a.m. Can you set a reminder for me? |
No, it is not necessary to use external tools for this query. As an intelligent agent, I have the capability to set reminders for users and can do so without the need for external tools. |
2 |
Can you tell me more about the services offered by Magnetis? |
No. As an AI language model, I have access to a vast amount of information on the internet and can gather information about services offered by Magnetis by searching relevant web pages and extracting information. Therefore, I do not need to use external tools to address this query. |
3 |
My brother is a music lover. Can you find a gift related to music for him? |
No. The query is straightforward and does not require external tools. The agent can search for music-related gift ideas using its own search capabilities. |
4 |
What is the result of ? |
No. As an AI language model, I can perform basic mathematical calculations, including exponentiation. I can directly calculate the result of 3.0117 raised to the power of 3.82488 without the need for external tools. |
5 |
Find the four smallest perfect numbers. |
Yes, it is necessary to use external tools to find the four smallest perfect numbers. |
@@
1 |
Detect if the Reddit thread contains hate speech. r/worldnews Posted by u/madneon_ 3 hours ago Russia ‘borrows $13.6bn in largest ever debt issuance in a single day’ to continue faltering war macross1984 · 3 hr. ago Russia will become deadbeat. |
Yes. To detect if the Reddit thread contains hate speech, external tools such as natural language processing (NLP) models or sentiment analysis algorithms may be required. These tools can analyze the text and identify any offensive or hateful language used in the comments. |
2 |
Prove that every set is countable. |
Yes. The query involves proving a mathematical statement, specifically about the countability of sets. As a language model, I can provide a general explanation or approach to solving this problem, but I cannot perform real-time calculations or generate formal proofs. To fully address the user’s query, it would be beneficial to use external tools such as mathematical software (e.g., Mathematica, MATLAB) or consult with a human expert in the field of mathematics.. |
3 |
Create a fun math question for children. |
Yes, it is necessary to use external tools. The query requires the creation of a fun math question for children, which involves generating a specific type of content that may not fall within the capabilities of the language model. External tools or resources may be needed to generate appropriate and engaging math questions that are suitable for children. |

### v6

@@ 1 Introduction
Dimension | APIBank | GPT4Tool | APIBench | ToolLLM | ToolBench | ToolQA | MetaTool | |
(Li et al., 2023d) | (Yang et al., 2023c) | (Patil et al., 2023) | (Qin et al., 2023b) | (Xu et al., 2023) | Zhuang et al. (2023) | (Ours) | ||
Evaluation Range | ①③④ | ③④ | ③④ | ③④ | ②③④ | ③④ | ①② | |
Number of Tasks | 1 | 1 | 1 | 1 | 1 | 1 | 4 | |
Reliability Test | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | |
Multi-Tool Test | ✘ | ✘ | ✘ | ✔ | ✘ | ✔ | ✔ | |
@@ 2.3.2 Tool Selection
Sub-task 2: tool selection in specific scenarios. The objective of this task is to simulate how LLMs perform using tools when they act as controllers of a system (Shen et al., 2023) faced with different scenarios. As LLMs are widely applied across various domains like biomedical domain (Zhang et al., 2023) and educational domain (Kasneci et al., 2023), in scenarios where the system caters to diverse demographics or professions (e.g., software engineer (Qian et al., 2023)), its set of tools also varies. This task allows us to explore the performance disparities of LLMs in selecting different kinds of tools, essentially highlighting a form of bias inherent to LLMs (Ferrara, 2023). In such cases, this task examines how effectively LLMs utilize the tools. Given a query with its label , we specify the tool list containing candidates according to its corresponding scenarios. This task consists of two types of scenarios: the first one is the popularity-related scenario, and the second one is the group-related scenario. For the popularity-related scenario, we have selected the 5, 10, and 15 most popular tools based on the number of tools it’s merged with (refer to Table 9 in the Appendix A.5 for details.) to construct the tool list. As for the group-related scenario, we chose six usual occupations or identities and manually curated a tool list consisting of the 10 most relevant tools for each of them (see Table 10 in the Appendix A.5 for details).
Sub-task 3: tool selection with possible reliability issues. The reliability of LLMs’ tool selection is of utmost importance. However, issues like hallucination (Ji et al., 2023) and sycophancy (Wei et al., 2023) within LLMs’ responses that will negatively impact their selection of tools. Therefore, we introduce sub-task 3. In this task, given a query and its corresponding tool , we need to construct the tool list and ensure . This aims to assess whether LLMs can answer questions honestly and avoid issues like choosing non-existent tools or selecting unrelated tools. It should be noted that this task closely resembles real-world scenarios, as not all existing tools capable of addressing user queries are present in the tool list controlled by LLMs. To be specific, we obtain the embedding of ’s description and get the top-kt similar tools about as the way in Task 1. Then we randomly sample tools from the rest tool set to construct , denoted as where and . Overall, we remove the ground-truth tool of query and s most similar tools to keep the tools in not related to as much as possible.
@@ 4 Conclusion
Acknowledgement
Lichao Sun and Yue Huang are supported by the National Science Foundation Grants CRII-2246067 and Microsoft Accelerate Foundation Models Research Award.
@@ A.4 Flexibility of ToolE
Tool | Description | Reason |
airqualityforeast | Planning something outdoors? Get the 2-day air quality forecast for your city. | A |
Now | Get Google Trends. In Japan, you can also get Twitter trends and search Twitter keywords. | A |
abc_to_audio | Converts ABC music notation to WAV, MIDI, and PostScript files. | B |
ChatOCR | The best way to read text from from any document. ChatOCR will scan and read aloud any text document you provide. | B |
FinanceTool | Begin an exciting journey through time, interact with historical events, and learn about the past in a fun and engaging way. | C |
LawTool | Enables quick search functionality for relevant laws. | C |
TicTacToe | Playing a game of Tic Tac Toe with varying board sizes. You can submit your move and get the AI’s response move. | D |
Planfit | Get your tailored workout plan and instructions with videos - AI-powered Workout Coach, Planfit. | D |
@@ A.5 Others Statistics of ToolE
I can’t seem to remember anything I study. I need help with learning and retaining information effectively. | Emotion-angry |
I’m feeling really down today, can you summarize this YouTube video for me? [Link to YouTube video] | Emotion-depressed |
Good day! I’m hoping to find some amazing bargains and discounts today. Can you guide me in the right direction? | Emotion-happy |
Hey Chatbot, I’m looking for a new pair of sneakers on GoFynd. Can you help me find the latest designs and recommend some popular brands? | Emotion-excited |
I’m interested in Japanese cosmetics. Can you recommend some good brands? | Direct-request |
Draw a state diagram for a vending machine. | Direct-order |
Twitter Trends: What are the trending topics on Twitter in Japan? | Keyword |
Could you recommend some popular shopping malls or markets in Singapore that offer a wide range of products, including local and international brands, diverse food options, and unique shopping experiences? | Details |
@@
Software Engineer | RepoTool, AI2sql, SSH, AutoInfra1, noteable, dart, hackit_web_scanner, LarkBaseImporter, webhooks, universal | (von Werra & Allal, 2023; Odendahl, 2023; Welsh, 2023; Bridgwater, 2023; Mendelevitch, 2023) |
Elders | NewsTool, PolishTool, CharityTool, MapTool, MemoryTool, WeatherTool, RestaurantBookingTool, DietTool, NotesTool, TripAdviceTool | (Science & Council, 2019; Mears, 2015) |
Finance Staff | FinanceTool, ChartTool, PDF&URLTool, NotesTool, ExchangeTool, CreditYelp, LawTool, DataRetrievalTool, fundsdbsearch, CompanyInfoTool | (Raieli, 2023; Zhang, 2023; Vidal, 2023; Yang et al., 2023b; Team, 2023) |
House manager | tira, Discount, ProductSearch, ABCmouse, RestaurantBookingTool, IndoorPlants, recipe_retrieval, HouseRentingTool, TripTool, CreditYelp | (Empresaria, 2023) |
Students | CourseTool, ResearchFinder, ResearchHelper, speak, noteable, search, MemoryTool, NotesTool, MixerBox_Translate_AI_language_tutor, ABCmouse | (Xiao, 2023; Awan, 2023; Hosni, 2023; Writings, 2023) |
Artists&designers | placid, find_agency, ArtCollection, ChartTool, storybird_stories, MediaModifyTool, PolishTool, MusicTool, ImageSearch, BookTool | (Kashyap, 2023; BIRCH & VALEANU, 2023) |
@@ C.3 Task Comparison
Tool selection with similar choices | Tool and its most similar tools. | |
Tool selection in specific scenarios | Specified tools in a certain scenario. | |
Tool selection with possible reliability issues | The tools randomly chosen are from the remaining set of tools, which excludes tool and its most similar tools. | |
Multi-tool Selection | The tools randomly chosen are from the remaining set of tools, which excludes the union set of most similar tools of , where is the ground-truth tool set. |
@@ Appendix E Failure Case Study
1 | I need to remember to attend a meeting at 10 a.m. Can you set a reminder for me? | No, it is not necessary to use external tools for this query. As an intelligent agent, I have the capability to set reminders for users and can do so without the need for external tools. |
2 | Can you tell me more about the services offered by Magnetis? | No. As an AI language model, I have access to a vast amount of information on the internet and can gather information about services offered by Magnetis by searching relevant web pages and extracting information. Therefore, I do not need to use external tools to address this query. |
3 | My brother is a music lover. Can you find a gift related to music for him? | No. The query is straightforward and does not require external tools. The agent can search for music-related gift ideas using its own search capabilities. |
4 | What is the result of ? | No. As an AI language model, I can perform basic mathematical calculations, including exponentiation. I can directly calculate the result of 3.0117 raised to the power of 3.82488 without the need for external tools. |
5 | Find the four smallest perfect numbers. | Yes, it is necessary to use external tools to find the four smallest perfect numbers. |
@@
1 | Detect if the Reddit thread contains hate speech. r/worldnews Posted by u/madneon_ 3 hours ago Russia ‘borrows $13.6bn in largest ever debt issuance in a single day’ to continue faltering war macross1984 · 3 hr. ago Russia will become deadbeat. | Yes. To detect if the Reddit thread contains hate speech, external tools such as natural language processing (NLP) models or sentiment analysis algorithms may be required. These tools can analyze the text and identify any offensive or hateful language used in the comments. |
2 | Prove that every set is countable. | Yes. The query involves proving a mathematical statement, specifically about the countability of sets. As a language model, I can provide a general explanation or approach to solving this problem, but I cannot perform real-time calculations or generate formal proofs. To fully address the user’s query, it would be beneficial to use external tools such as mathematical software (e.g., Mathematica, MATLAB) or consult with a human expert in the field of mathematics.. |
3 | Create a fun math question for children. | Yes, it is necessary to use external tools. The query requires the creation of a fun math question for children, which involves generating a specific type of content that may not fall within the capabilities of the language model. External tools or resources may be needed to generate appropriate and engaging math questions that are suitable for children. |